# Mining the time axis with TRON. I. Millisecond pulsars in Omega Centauri, Terzan 5 and 47 Tucanae detected through MeerKAT interferometric imaging

### v1

@@ 1 Introduction
MeerKAT’s high instantaneous sensitivity and large field of view make it an excellent instrument for detecting radio transients. The ThunderKAT large science program (LSP) has yielded a wide range of important discoveries (see e.g. Andersson et al., 2022, 2023; Driessen et al., 2022, 2020). Typically, these have been either triggered by dedicated transient search engines such as MeerTRAP (Rajwade et al., 2020; Malenta et al., 2020), or have involved targeted multi-epoch observations of particular fields. Recent results (e.g. chastain2023; dobie2023; Heywood, 2023; wang2023; deruiter2024; fijma2024; Smirnov et al., 2024a) suggest that more routine imaging observations can also yield a rich vein of medium-timescale (minute-to-hour) radio transients or variables. For example, Smirnov et al. (2024a) report on the serendipitous discovery of the PARROT, a pulsar-type object, during observations of the Great Saturn–Jupiter Conjunction of 2020, while Heywood (2023) reports a millisecond pulsar (MSP) candidate detected in MeerKAT imaging observations of the globular cluster 47 Tucanae (47 Tuc). In both cases, medium-timescale spectro-temporal variability (whether intrinsic, or caused by refractive effects along the line-of-sight) is the observational signature. With almost six years’ worth of MeerKAT imaging observations already in the archive, a dedicated program of mining this data for more such discoveries seems well worth the effort.
@@ 2 Observations and data reduction
  • Smoothing the raw time resolution cube to a set of longer timescales so as to increase sensitivity to longer-duration events (the timescales are a tunable parameter, e.g. 15 s, 30 s, 60 s, etc.).
@@ 4 Conclusions
We detect two eclipsing spider MSPs, two highly scintillating MSPs, and one candidate MSP showing hints of eclipsing. There are 12 known eclipsing spiders within these three fields; a follow-up study to investigate why the other eclipsing systems are not being detected via this method would be interesting, and could serve to refine the TRON approach. The current version of TRON only utilizes wideband HTC images and lightcurves, which must tend to wash out chromatic effects such as scintillation. Expanding TRON to a more fine-grained search of the frequency axis would be a very promising development. In the meantime, it is already clear that eclipsing and scintillation are observational signatures that could be used to uncover more pulsars via image-plane search techniques.

### v2

@@ 1 Introduction
MeerKAT’s high instantaneous sensitivity and large field of view make it an excellent instrument for detecting radio transients. The ThunderKAT large science program (LSP) has yielded a wide range of important discoveries (see e.g. Andersson et al., 2022, 2023; Driessen et al., 2022, 2020). Typically, these have been either triggered by dedicated transient search engines such as MeerTRAP (Rajwade et al., 2020; Malenta et al., 2020), or have involved targeted multi-epoch observations of particular fields. Recent results (e.g. Chastain et al., 2023; Dobie et al., 2023; Wang et al., 2023; de Ruiter et al., 2024; Fijma et al., 2024) suggest that more routine imaging observations can also yield a rich vein of medium-timescale (minute-to-hour) radio transients or variables. For example, Smirnov et al. (2024a) report on the serendipitous discovery of the PARROT, a pulsar-type object, during observations of the Great Saturn–Jupiter Conjunction of 2020, while Heywood (2023) reports a millisecond pulsar (MSP) candidate detected in MeerKAT imaging observations of the globular cluster 47 Tucanae (47 Tuc). In both cases, medium-timescale spectro-temporal variability (whether intrinsic, or caused by refractive effects along the line-of-sight) is the observational signature. With almost six years’ worth of MeerKAT imaging observations already in the archive, a dedicated program of mining this data for more such discoveries seems well worth the effort.
@@ 2 Observations and data reduction
  • Smoothing the raw time resolution cube to a set of longer timescales so as to increase sensitivity to longer-duration events. The set of timescales is a tunable parameter, with the default setting of 15 s to 960 s in successive increments of ×$\times 2$ used in this work.
@@
    For this work, we used the current default Breifast settings of a $7\sigma$ peak threshold (decreasing to $6\sigma$ at the longest two timescales) with respect to a local r.m.s. estimate, and a minimum 10% excursion relative to the baseline flux measured in the deep image. This served to exclude any false-positives that would otherwise be induced by instrumental effects such as primary beam rotation and varying $uv$-coverage.
@@ 4 Conclusions
We detect two eclipsing spider MSPs, two highly scintillating MSPs, and one candidate MSP showing hints of eclipsing. There are 12 known eclipsing spiders within these three fields; a follow-up study to investigate why the other spider systems are not being detected via this method will be used to refine the TRON approach. The current version of TRON only utilizes wideband HTC images and lightcurves, which must tend to wash out chromatic effects such as scintillation. Expanding TRON to a more fine-grained search of the frequency axis would be a very promising development. In the meantime, it is already clear that eclipsing and scintillation are observational signatures that could be used to uncover more pulsars via image-plane search techniques.